# Density of states and non-smooth Lyapunov exponent in the localized phase

### v1

@@ Abstract
Conclusion and Discussion.—The localization of wave function in the disordered potential can be characterized by the Lyapunov exponent $\gamma(E)$, in which the transition from extended state to the localized state is marked by $\gamma(E)$ changes from zero to finite value [62, 63, 64, 65]. This is a continuous function of $E$, yet it is not smooth. In this work, we address the fundamental question that whether $\gamma(E)$ can be non-smooth in the disordered models, which is not prohibited by its definition. We present a simple intuitive mechanism that when the system has two (or more) localized states, it is possible to observe non-smooth behavior of $\gamma(E)$. We considered the one dimensional model with slowly varying potential [45, 46, 47, 48, 49], which support localization at the potential extrema and potential nodes, and show that in these particular models the non-smoothness of this function is attributed to the singularities of the DOS [44]. In this work we tentatively consider all localized states to be the same phase, and in future, it is quite possible that if new classification criteria are invented, these states may also be categorized into different phases. Since this non-smoothness of the Lyapunov exponent is in principle allowed in theory as demonstrated in this manuscript, this idea may be generalized to higher dimensions, which should greatly broaden our understanding of Anderson localization [1, 2, 3, 4, 5, 6, 7, 8, 9, 66].
@@
We thank Qi Zhou for valuable discussion. This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0500000), the National Natural Science Foundation of China (Grant No. 12374017, No. 12074362) and the Innovation Program for Quantum Science and Technology (2021ZD0303303, 2021ZD0301200, 2021ZD0301500).

### v2

@@ IV Conclusions
The localization of wave functions in disordered potentials can be characterized by the Lyapunov exponent $\gamma(E)$, whose transition from zero to a finite value marks the change from extended to localized states [54, 55, 56]. Although $\gamma(E)$ varies continuously with energy, it is generally non-analytic across localization transitions. In this work, we address the fundamental question of whether $\gamma(E)$ can be non-analytic in the localized phase, which is not prohibited by its definition. We propose an intuitive mechanism for this phenomenon that when the system supports two (or more) kinds of localized states, the Lyapunov exponent can display non-analytic behavior within the localized phase. To illustrate this idea, we analyzed one-dimensional models with slowly varying potential [39, 40, 41, 42, 43], where we observe a non-analytic Lyapunov exponent in the localized phase. Further investigation shows that these systems support two kinds of localized states which localize at the potential extrema and potential nodes, and the boundaries for these localized states are the transition points of the non-analytic Lyapunov exponent. According to the Thouless formula, the non-analyticity of the Lyapunov exponent is attributed to the singularities of the DOS [38]. Finally, we extend our discussion to one-dimensional models with slowly varying hopping and find similar phenomena. In this work we tentatively consider all localized states to be the same phase, and in future, it is quite possible that if new classification criteria are invented, these states may also be categorized into different phases. Since this non-analyticity of the Lyapunov exponent is in principle allowed in theory as demonstrated in this manuscript, this idea may be generalized to higher dimensions, which should greatly broaden our understanding of Anderson localization [1, 2, 3].
Acknowledgements.
@@
We thank Profs. Zhenyu Zhang, Qian Niu, and Qi Zhou for valuable discussion. This work is supported by the National Natural Science Foundation of China (Grant No. 12374017, No. 12074362 and No. U23A2074), the Innovation Program for Quantum Science and Technology (2021ZD0303303, 2021ZD0301200, 2021ZD0301500) and the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0500000).